If tan 3x = sin45*cos45+sin30 then find the value of x.( with steps)

Grade:9

1 Answers

Soumendu Majumdar
5 years ago
Dear Student,
Taking R.H.S we get,
1/2 + 1/2 =1   since sin 45^{\circ} = cos 45^{\circ} = 1/\sqrt{2}   and sin 30^{\circ} =1/2
Now 
tan 3x =1 = tan 45^{\circ}
so x = 15^{\circ}
Hope it helps!
 
with regards,
Soumendu
P.s- If you have any confusion regarding the solution feel free to ask, I’ll be glad to help you!